OpenStudy (jpsmarinho):

Lim x* sen(1/x) when x goes to 0. I did and obtain 1, but, in the book, the answer is 0!

OpenStudy (anonymous):

Did you use a substitution?

OpenStudy (anonymous):

Just keep in mind, if you perform the following substitution \[ u = \frac{1}{x} \] you want to change the limit that is being approached, in order to get the correct answer. When x approaches infinity, u goes to zero. Otherwise, use a Series representation.

OpenStudy (jpsmarinho):

Which is Series representation?

OpenStudy (jpsmarinho):

Do this is wrong? \[\lim_{x \to 0} \frac{sin(1/x)}{1/x}=1\]

OpenStudy (anonymous):

0 is the answer

OpenStudy (anonymous):

that is not the answer. :( You can use the squeeze thm for this.

OpenStudy (anonymous):

@rainaby .....Don't just give answers man.

OpenStudy (anonymous):

no its simple see for any x sin(x) is finite and multiply it with 0 u get 0 as answer

OpenStudy (anonymous):